OpenStudy (anonymous):

What is 5.8 × 106 mg/s in grams per minute? A. 9.7 × 109 g/min B. 9.7 × 107 g/min C. 3.48 × 105 g/min D. 3.48 × 103 g/min

OpenStudy (dangerousjesse):

\(36.89~~ g/min\)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

So what's the answer?

OpenStudy (tylerd):

should be 36.888 grams per minute

OpenStudy (anonymous):

But that's not an option

OpenStudy (tylerd):

yep so either they are wrong, or you miss typed the answeres given here.

OpenStudy (anonymous):

Oh i forgot to input the exponents

OpenStudy (aum):

5.8 × 10^6 mg/s * ( 1 gm / 1000 mg ) / (1/60 min / sec) = 5.8 × 10^6 x 10^(-3) * 60 = ? gm / min

OpenStudy (dangerousjesse):

Yeah, I normally assume teachers want you to round to the nearest hundredth, sorry.

OpenStudy (aum):

3.48 × 10^5 g/min

OpenStudy (anonymous):

A. \[9.7 × 10^{9} g/\min\] B. \[9.7 × 10^{7} g/\min\] C. \[3.48 × 10^{5} g/\min\] D. \[3.48 × 10^{3} g/\min\]

OpenStudy (anonymous):

those were the answers

OpenStudy (aum):

5.8 × 10^6 x 10^(-3) * 60 = 3.48 × 10^5 g / min

OpenStudy (anonymous):

so C?

OpenStudy (anonymous):

Yes, C. Aum explained it.